Suppose that a change in the price of bread from R10 to R12.50 leads to a change in the quantity demanded of bread from 65 units to 90 units. Use the arc elasticity method to calculate and interpret the price elasticity of demand for bread
1
Expert's answer
2020-10-12T10:29:03-0400

To calculate arc elasticity of demand , we first take the midpoint in between.

Midpoint "Q=\frac{(Q1+Q2)}{2}"


Midpoint P="\frac{(P1+P2)}{2}"

Therefore;

Arc elasticity of demand (PED)= "\frac{\frac{Q2-Q1}{Q}}{\frac{P2-P1}{P}}"


Mid point of Q="\frac{65+90}{2}" =77.5


Mid point of P="\frac{(10+12.5)}2=11.25"


% change in Q ="\frac{(90-65)}{77.5}=0.3226"


% change in P="\frac{12.50-10}{11.25}=0.2222"

PED="\frac{0.3226}{0.2222}=1.45"